OpenStudy (anonymous):

Find the average rate of change of g(x)=-x^2+2x+2 from x=2 to x=6

jhonyy9 (jhonyy9):

for x=2 you get 10 for x=6 ---- 50 what is the average rate ?

OpenStudy (anonymous):

just consider the end points, it doesn't matter what the function does in between. g(6)-g(2)/(6-2)

OpenStudy (anonymous):

would it be 4?

OpenStudy (whpalmer4):

average rate of change of a function \(f(x)\) between the points \((a,f(a))\) and \((b,f(b))\) is simply the slope of a line through those two points, and therefore is \[\frac{f(b)-f(a)}{b-a}\]

OpenStudy (whpalmer4):

what is \(g(6)=\) what is \(g(2)=\)

OpenStudy (anonymous):

g(6)= -22 and g(2)= 2

OpenStudy (whpalmer4):

okay, what is \[\frac{g(6)-g(2)}{6-2}=\]

OpenStudy (anonymous):

-24/4 ?

OpenStudy (whpalmer4):

Here's a graph showing both the curve (in blue) and the line passing through those two points.

OpenStudy (whpalmer4):

yes, slope is \(-24/4 = -6\)

OpenStudy (anonymous):

So -6 is the average rate of change?

OpenStudy (whpalmer4):

yes. the average rate of change is just the slope of the line that goes through the endpoints of the region of interest. matters not a whit what the function does in between them.